Assume that GDP is $6,000, personal disposable income is $5,100, and the government bud


get deficit is $200. Consumption is $3,800, and the trade deficit is $100.


a. How large is saving (S)? b. How large is investment (1)?


c. How large is government spending (G)?


1
Expert's answer
2021-08-12T05:09:02-0400

a. The disposable income is either consumed or saved.

"YD=C+S \\
S=YD-C \\
= 5100-3800 \\
= 1300"

The saving is $1300.

b. To calculate the investment, use the relation between saving, investment, government budget and trade.

"S-I = (G+TR-TA)+NX \\

I = (TA-TR-G) -NX +S \\

= -200 -100 +1300 \\

= 1200"

The investment is $1200.

c. To calculate the government spending, use the national income identity.

"Y=C+I+G+NX \\

G= Y-C-I-NX \\

= 6000 -3800- 1200 -(-100) \\

= 1100"

The government spending is $1100.